OpenStudy (anonymous):

what is the derivative of (x^2+1)^(arcsin(x))?

OpenStudy (anonymous):

having trouble on what to do.

OpenStudy (anonymous):

what do you mean by taking the log

OpenStudy (anonymous):

for the second method i got e^((x2+1)ln(arcsin(x))) and then the derivative of (x2+1)ln(arcsin(x)

OpenStudy (anonymous):

i dont understand how to do the first method however

OpenStudy (anonymous):

wait how do you take the derivative of a log though

OpenStudy (anonymous):

it is a product rule and also a chain rule problem

OpenStudy (anonymous):

ehhh you're talking about the second method right?

OpenStudy (anonymous):

the derivative of \(\ln(f(x))\) is \(\frac{f'(x)}{f(x)}\) either method, it amounts to exactly the same work

OpenStudy (anonymous):

when you're saying you take the log of both sides you mean ln(y)=ln(x^2+1)^(arcsinx) right and not log(y)=log(...)

OpenStudy (anonymous):

i know how to take the derivative of ln(x) its 1/x times x'

OpenStudy (anonymous):

it is \[\ln(y)=\] oh damn i had it backwards!!

OpenStudy (anonymous):

\[y=(x^2+1)^{\arcsin(x)}\] \[\ln(y)=\arcsin(x)\ln(x^2+1)\] thats better, sorry

OpenStudy (anonymous):

now take the derivative using the chain rule and product rule

OpenStudy (anonymous):

ok so i get 1/y = dy/dx arc sin(x) (1/x^2+1)(2x)

OpenStudy (anonymous):

and i divide everything by arc sin(x)(1/x^2+1)(2x) to have dy/dx on one side and i get arc sin(x)(1/x^2+1)(2x)/y = dy/dx?

OpenStudy (anonymous):

\[\ln(y)=\arcsin(x)\ln(x^2+1)\] \[\frac{y'}{y}=\frac{1}{\sqrt{1-x^2}}\ln(x^2+1)+\arcsin(x)\frac{2x}{x^2+1}\]

OpenStudy (anonymous):

product rule and chain rule

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

yw then to finish multiply all this mess by the original function